# Least Squares with Equality constraints Extreme Learning Machines for the resolution of PDEs

### v2

@@ 2 Introduction
The starting points for the study of the present paper are PIELM, as named in [22], and their alternative named XTFC in [49]. PIELM, as already pointed out in [5], is a collocation method that is eventually solved in the LS sense. XTFC adds to this method the use of Functional Connections to treat the boundary conditions analytically. Thanks to this, many problems can be efficiently solved, especially in the 1D setting (ordinary differential problems, two points boundary problems, optimal control) or with tensor product domains but equivalent approaches for general domains are still lacking,. In the present paper, we propose a different approach to overcome this limitation and give insights concerning the notion of convergence in such a setting. The last is an important theoretical and practical question that was already addressed in [10] in the case of interpolation, whereas generally the focus on the use of NN for the resolution of PDEs only focuses on accuracy. The question well arises in the construction with ELM because no iteration is needed for the training as the overall training problem is linear. In the present paper, the convergence is discussed when the training set is increased (more collocation points); when the number of degrees of freedom is increased (more neurons in the hidden layer); and finally when both are increased keeping fixed the ratio between the two.
This gives that we have the possibility to obtain accurate solutions that in many PINN strategies are not obtained, due to the fixed number of degrees of freedom and the suboptimality of nonconvex nonlinear optimizations. The structure of ELMs, which in our personal vision bridge spectral methods and ”pure” PINNs, allows an easy definition of spectral convergence. In our numerical tests, we can also see nice convergence and high accuracy in challenging problems.
@@
Now, let us introduce some preliminary notation. In the rest of the paper we will focus on linear elliptic problems of the type
@@
where $\Omega\subset\mathbb{R}^{2}$ is an arbitrary domain and $\Gamma$ is a curve; usually $\Gamma=\partial\Omega$.
@@
The parameters $\bm{\beta}\in\mathbb{R}^{L}$ are the unknowns, determined via some exactness condition imposed on a set of $N$ points, where in general we have $N_{I}$ of them inside $\Omega$ and $N_{B}$ on the curve $\Gamma$.
@@ 2.2 The use of ELMs for the resolution of PDEs: Review of the literature
The first paper that uses a method that relays on ELM for the resolution of PDEs is [51]. In this paper the ELM are not considered as the ansatz functions. The first articles that introduce the use of ELM networks as ansatz functions in the resolution of linear PDEs are [5, 22], dealing with both linear problems and the soft imposition of Dirichelet boundary conditions. In particular, [5] focuses on elliptic problems with boundary layers, while [22] results in 2D general-shape domains and discusses the accuracy and versatility of the procedure. Then, in the same year Schiassi et al. introduce the XTFC in [49] and test accuracy when the boundary condition imposed hardly. XTFC are then used in various PDE applications, see [49, 47, 46, 54].
Starting from these results, ELMs have been used in different applications. Linear elasticity problems are considered in [56, 59]. The Stokes equations is solved in [6]. Higher order PDEs problems have been considered in [23, 33]. Nonlinear PDEs have been considered firstly in [18, 21, 26] and more recently in [29, 37, 52]. Inverse problems also are considered in [19].
@@ 3.1 ELM Neural Networks
In this paper we focus on linear scalar equations in two dimensions, while reporting theoretical results, we emphasize that such results could be reported in a more general framework. A Single Layer Feedforward Neural Network (SLFN) with $L$ hidden neurons is determined by the input weights $\{\bm{w}_{j}\in\mathbb{R}^{2}\}_{1\,\leq\,j\,\leq L}$, the hidden neurons’ biases $\{b_{j}\in\mathbb{R}\}_{1\,\leq\,j\,\leq L}$, the output weights $\{\beta_{j}\in\mathbb{R}\}_{1\,\leq\,j\,\leq L}$, and an activation function $\sigma:\mathbb{R}\to\mathbb{R}$. In what follows, we restrict our attention to the case in which $\sigma$ is the hyperbolic tangent and the $\bm{w}_{j}$ and $b_{j}$ are sampled independently from the uniform distribution on $[-M,M]$, where the parameter $M$ controls how steep are the resulting basis functions.
In order to fix the notation for the training of the SLFN, let us consider the following approximation problem, see [10]. Given $N$ samples $(\bm{x}_{i},y_{i})\in\mathbb{R}^{2}\times\mathbb{R}$, the SLFN aims to approximate the data $y_{1},\ldots,y_{N}$ with weighted sums; in formula:
@@
ELM is a training algorithm for SLFNs where the input weights $\{\bm{w}_{j}\}_{{1\leq j\leq n}}$ and the biases of hidden neurons $\{b_{j}\}_{{1\leq j\leq n}}$ are randomly selected from a continuous distribution and remain fixed, the optimization is performed exclusively on the output weights vector. Since the training set is also fixed, the entries of the matrix
@@ 3.3 Imposition of boundary conditions
    $\begin{cases}\mathcal{D}[u_{I}](\bm{x})=f(\bm{x})-\mathcal{D}[u_{B}](\bm{x})\qquad&\bm{x}\in\Omega\\ \mathcal{B}[u_{I}](\bm{x})=0\qquad&\bm{x}\in\partial\Omega\end{cases}$ (11)
@@
We can introduce $\bm{A}\in\mathbb{R}^{N_{I}\times L}$ and $\bm{f}\in\mathbb{R}^{N_{I}}$ such that:
@@
The previous strategy can be generalized to convex polynomials [43], and using different interpolation strategies it is possible to work on arbitrary polygons. Strategies that allow to work on arbitrary domains [24] exist but unfortunately, analytical expressions for the interpolants are not available and their evaluations lead to quadratures or numerical optimization problems that are harder and require more computations than the original problem. On top of all of these complications, everything scales up with the dimensions of the problem [27].
We are now interested in the introduction to soft imposition of boundary conditions, that is probably the most common approach in the PINN literature, as it asks only to perform collocation on the boundary as well. This has also been used for ELMs as Physics-Informed ELM (PIELM) in [22]: the collocation points $\{\bm{x}_{k}\}_{{1\leq k\leq N_{B}}}$ on $\Gamma$ are added to the training set and we define $\bm{C}\in\mathbb{R}^{N_{B}\times L}$ and $\bm{g}\in\mathbb{R}^{N_{B}}$ such that:
@@
Analogously we define $\bm{A}\in\mathbb{R}^{N_{I}\times L}$ and $\bm{f}\in\mathbb{R}^{N_{I}}$ such that:
@@
And the collocation problem is the following:
$\bm{\beta}^{*}=\arg\min\Bigl{(}\frac{1}{2}\|\bm{A\beta}-\bm{f}\|_{2}^{2}+\frac{\lambda}{2}\|\bm{C\beta}-\bm{f}\|_{2}^{2}\Bigr{)}$ (18)
@@ 4 Comparisons
We fix the problem 24 so to have a solution with high regularity. In particular, we take, see Figure 3:
@@
As a general comment, we can notice that the accuracy for the three procedures is similar, but some remarks on the computational costs can be highlighted. PIELM solves a unique linear system of dimension $N\times L$, XTFC also solves a unique system of the same type, but we remark that in this case $N=N_{I}$ and a pre-processing for the resolution of the boundary problem is needed. The proposed procedure, on the other hand, solves two linear systems, one of dimension $N_{B}\times L$ and the final one with $N_{I}$ rows COLONNE. This gives a saving in terms of computational cost and has an impact on the condition number of the overall resolution. Following what was done in [Chen2024], we present in Figure 6 the comparison between the spectra obtained via SVD. Fixing the dimension $N=100,300,3000$, in the plots we present both the square and the $L=2N$ overparameterized case. Different colors refer to different matrices: the collocation of the original ELM functions on the $N_{I}$ internal points (denoted by ”original”), the collocation on all $N_{I}+N_{B}$ points for the PIELM method, the collocation of the modified functions on the $N_{I}$ internal points for the LSE-ELM method, the collocation on $N$ (all internal) points for the XTFC method. From the plot, it appears that the spectra reach machine precision after $\approx 250$ values, we remark that this could depend on the very simple structure both of the problem and of the domain.
@@ 5.1 L-shape and star shape domain
The Poisson problem is solved in a square domain and the fixed conditions are imposed on $\Gamma$ that thus not coincides with the boundary. On the top, the values are enforced on a star inside the domain. On the bottom, the values are enforced on an L-shaped region that partially intersect with the boundary of the domain. In these tests we use $N_{I}=2789,N_{B}=211$. In both cases we obtain very accurate results inside and good results also in the external part.
@@ 6 Conclusions
In the present work, we have presented a new method for the numerical resolution of differential problems using single hidden-layer neural networks. This strategy is able to consider boundary conditions in an efficient way and mitigates some negative aspects of the previous strategies. The proposed LSE-ELM method is a global collocation strategy, and this gives many pros and some contras. The pros include the possibility to consider points outside the domain in the resolution, so as to move the computation in the direction of physically based extrapolation. In fact, in the resolution of PDEs from the point of view of training, one can consider that the usual boundary conditions are considered good ”supervised data”, while the satisfaction of the differential problem can be considered a ”physical constraint”. Then, the resolution both inside the domain and outside can be an important feature in applications where extrapolation is required. The contras include the resolution of a full matrix and the loss of accuracy when the problem is less regular. The latter is an issue that can be fixed only with local strategies or the use of ad hoc activation functions. For the first contra, we have noticed with some computations that the spectra are better than the direct competitors, and we believe that some adaptation can further improve such a task. The constructions in the present paper are done in the most general case, and no adaptivity has been included, as an example with the use of Null Rules [2, 3] one could adapt the shape of the functions. However, the strategy has been used in such a setting in order to emphasize the ability of the method to deal with randomly chosen points and activation functions.

### v3

@@ 2 Introduction
The starting points for the study of the present paper are PIELM, as named in [23], and their alternative named XTFC in [51]. PIELM, as already pointed out in [5], is a collocation method that is eventually solved in the LS sense. XTFC adds to this method the use of Functional Connections to treat the boundary conditions analytically. Thanks to this, many problems can be efficiently solved, especially in the 1D setting (ordinary differential problems, two points boundary problems, optimal control) or with tensor product domains but equivalent approaches for general domains are still lacking. In the present paper, we propose a different approach to overcome this limitation and give insights concerning the notion of convergence in such a setting. The last is an important theoretical and practical question that was already addressed in [11] in the case of interpolation, whereas generally the focus on the use of NN for the resolution of PDEs only focuses on accuracy. The question well arises in the construction with ELM because no iteration is needed for the training as the overall training problem is linear. In the present paper, the convergence is discussed when the training set is increased (more collocation points); when the number of degrees of freedom is increased (more neurons in the hidden layer); and finally when both are increased keeping fixed the ratio between the two.
This gives that we have the possibility to obtain accurate solutions that in many PINN strategies are not obtained, due to the fixed number of degrees of freedom and the suboptimality of nonconvex nonlinear optimizations. The structure of ELMs, which in our personal vision bridge spectral methods and ”pure” PINNs, allows an easy definition of spectral convergence. In our numerical tests, we can also see nice convergence and high accuracy in challenging problems, both linear and nonlinear, both stationary and time-dependent. The method is also tested in cases with minor regularity, and in this case, convergence is not guaranteed.
@@
Now, let us introduce some preliminary notation. In the rest of the paper we will focus on problems of the type
@@
where $\mathcal{D}$ is a differential operator (linear or nonlinear), $\mathcal{B}$ is a linear operator, $\Omega\subset\mathbb{R}^{d}$ is an arbitrary domain with $d=2,3$ and $\Gamma\subset\mathbb{R}^{d-1}$; usually $\Gamma=\partial\Omega$. In the rest, we will not distinguish between spatial and temporal variables, since the whole treatment will be carried out in the same way. Namely, we consider space-time discretizations and treat the initial as part of the boundary conditions, so that in these cases $\Gamma$ wil not include points at final time.
@@
The parameters $\bm{\beta}\in\mathbb{R}^{L}$ are the unknowns, determined via some exactness condition imposed on a set of $N$ points, where in general we have $N_{I}$ of them inside $\Omega$ and $N_{B}$ on $\Gamma$.
@@ 2.2 The use of ELMs for the resolution of PDEs: Review of the literature
The first paper that uses a method that relies on ELM for the resolution of PDEs is [53]. In this paper the ELM are not considered as the ansatz functions. The first articles that introduce the use of ELM networks as ansatz functions in the resolution of linear PDEs are [5, 23], dealing with both linear problems and the soft imposition of Dirichelet boundary conditions. In particular, [5] focuses on elliptic problems with boundary layers, while [23] results in 2D general-shape domains and discusses the accuracy and versatility of the procedure. Then, in the same year Schiassi et al. introduce the XTFC in [51] and test accuracy when the boundary condition imposed hardly. XTFC are then used in various PDE applications, see [51, 50, 48, 57].
Starting from these results, ELMs have been used in different applications. Linear elasticity problems are considered in [59, 62]. The Stokes equations are solved in [6]. Higher order PDEs problems have been considered in [24, 34, 54]. Nonlinear PDEs have been considered firstly in [19, 22, 27] and more recently in [30, 38, 55]. The moving boundaries for the Stefan problem are considered in [41]. Inverse problems are considered in [20].
@@ 3.1 ELM Neural Networks
In this paper we focus on linear scalar equations in two dimensions, while reporting theoretical results, we emphasize that such results could be reported in a more general framework. A Single Layer Feedforward Neural Network (SLFN) with $L$ hidden neurons is determined by the input weights $\{\bm{w}_{j}\in\mathbb{R}^{d}\}_{1\,\leq\,j\,\leq L}$, the hidden neurons’ biases $\{b_{j}\in\mathbb{R}\}_{1\,\leq\,j\,\leq L}$, the output weights $\{\beta_{j}\in\mathbb{R}\}_{1\,\leq\,j\,\leq L}$, and an activation function $\sigma:\mathbb{R}\to\mathbb{R}$. In what follows, we restrict our attention to the case in which $\sigma$ is the hyperbolic tangent and the $\bm{w}_{j}$ and $b_{j}$ are sampled independently from the uniform distribution on $[-M,M]$, where the parameter $M$ controls how steep are the resulting basis functions.
In order to fix the notation for the training of the SLFN, let us consider the following approximation problem, see [11]. Given $N$ samples $(\bm{x}_{i},y_{i})\in\mathbb{R}^{d}\times\mathbb{R}$, the SLFN aims to approximate the data $y_{1},\ldots,y_{N}$ with weighted sums; in formula:
@@
ELM is a training algorithm for SLFNs where the input weights $\{\bm{w}_{j}\}_{{1\leq j\leq L}}$ and the biases of hidden neurons $\{b_{j}\}_{{1\leq j\leq L}}$ are randomly selected from a continuous distribution and remain fixed, the optimization is performed exclusively on the output weights vector. Since the training set is also fixed, the entries of the matrix
@@ 3.2 Discretization of the operator: collocation
In the case of a nonlinear operator the training is a nonlinear optimization with the following residual:
$\mathcal{R}_{i}(\bm{\beta})=\mathcal{D}[\tilde{u}](x_{i})-f(x_{i})\qquad 1\leq i\leq N_{I}$ (11)
@@ 3.3 Imposition of boundary conditions
    $\begin{cases}\mathcal{D}[u_{I}+u_{B}](\bm{x})=f(\bm{x})\qquad&\bm{x}\in\Omega\\ \mathcal{B}[u_{I}](\bm{x})=0\qquad&\bm{x}\in\partial\Omega\end{cases}$ (12)
@@
For linear operators we can introduce $\bm{A}\in\mathbb{R}^{N_{I}\times L}$ and $\bm{f}\in\mathbb{R}^{N_{I}}$ such that:
@@
As already mentioned, the case of a nonlinear operator leads to a nonlinear optimization problem:
$\mathcal{R}_{i}(\bm{\beta})$ $=\mathcal{D}[u_{I}+u_{B}](\bm{x}_{i})-f(\bm{x}_{i})\qquad 1\leq i\leq N_{I}$
[…]
@@
The previous strategy can be generalized to convex polygons [45], and using different interpolation strategies it is possible to work on arbitrary polygons. Strategies that allow to work on arbitrary domains [25] exist but unfortunately, analytical expressions for the interpolants are not available and their evaluations lead to quadratures or numerical optimization problems that are harder and require more computations than the original problem. On top of all of these complications, everything scales up with the dimensions of the problem [28].
We are now interested in the introduction to soft imposition of boundary conditions, that is probably the most common approach in the PINN literature, as it asks only to perform collocation on the boundary. This has also been used for ELMs as Physics-Informed ELM (PIELM) in [23]: the collocation points $\{\bm{x}_{k}\}_{{1\leq k\leq N_{B}}}$ on $\Gamma$ are added to the training set and we define $\bm{C}\in\mathbb{R}^{N_{B}\times L}$ and $\bm{g}\in\mathbb{R}^{N_{B}}$ such that:
@@
The collocation for linear and nonlinear problems is as follows:
$\bm{\beta}^{*}$ $=\arg\min\Bigl{(}\frac{1}{2}\|\bm{A\beta}-\bm{f}\|_{2}^{2}+\frac{\lambda}{2}\|\bm{C\beta}-\bm{g}\|_{2}^{2}\Bigr{)}$ (18)
@@ 3.4 Constrained imposition of boundary conditions: the proposed method
This approach can be extended to nonlinear problems as follows:
  1. 1.
    Consider any nonlinear optimization algorithm that uses the Jacobian of the loss function
  2. 2.
    At each step, compute the Jacobian with respect to $\bm{z}$ by the chain rule:
[…]
  3. 3.
    Compute the update $\bm{\delta z}$ according to the chosen algorithm
  4. 4.
    Compute the updated set of parameters from $\bm{\beta+\delta\beta}$ where:
    $\bm{\delta\beta}=\bm{V}_{\perp}\bm{\delta z}$ (26)
@@ 4 Comparisons
We fix the problem 27 so to have a solution with high regularity. In particular, we take, see Figure 3:
@@
As a general comment, we can notice that the accuracy for the three procedures is similar, but some remarks on the computational costs can be highlighted. PIELM solves a unique linear system of dimension $N\times L$, XTFC also solves a unique system of the same type, but we remark that in this case $N=N_{I}$ and a pre-processing for the resolution of the boundary problem is needed. The proposed procedure, on the other hand, solves two linear systems, one of dimension $N_{B}\times L$ and the final one with $N_{I}$ rows. This gives a saving in terms of computational cost and has an impact on the condition number of the overall resolution. Following what was done in [7], we present in Figure 6 the comparison between the spectra obtained via SVD. Fixing the dimension $N=100,300,3000$, in the plots we present both the square and the $L=2N$ overparameterized case. Different colors refer to different matrices: the collocation of the original ELM functions on the $N_{I}$ internal points (denoted by ”original”), the collocation on all $N_{I}+N_{B}$ points for the PIELM method, the collocation of the modified functions on the $N_{I}$ internal points for the LSE-ELM method, the collocation on $N$ (all internal) points for the XTFC method. From the plot, it appears that the spectra reach machine precision after $\approx 250$ values, we remark that this could depend on the very simple structure both of the problem and of the domain.
@@ 5 Numerical tests
Finaly, we consider some examples of nonlinear problems, following the approach for the resolution of the nonlinearity presented in section 3.4. In particular two time-dependent nonlinear parabolic problems are used.
@@
In the case of time-dependent problems, the inial condition is teated as a fixed boundary condition, while no condition is imposed at the boundary of the domain corresponding to the final time.
@@ 5.1 L-shape and star shape domain
The Poisson problem is solved in a square domain and the fixed conditions are imposed on a curve $\Gamma$ that does not coincides with the boundary. On the top, the values are enforced on a star inside the domain. On the bottom, the values are enforced on an L-shaped region that partially intersect with the boundary of the domain. In these tests we use $N_{I}=2789,N_{B}=211$. In both cases we obtain very accurate results inside and good results also in the external part.
@@ 5.5 Reentrant corner
5.6 1D Burgers-Fisher equation
In this first time-dependent example, we focus on a problem with nonlinearities in the transport and reaction terms:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\frac{\partial^{2}u}{\partial x^{2}}+u(1-u)\ .$ (30)
The spatial domain is $\Omega=[0,5]$ and the equation is solved for values of $t$ in the interval $[0,5]$, resulting in a square domain for a space-time collocation approach. The Dirichlet boundary conditions are fixed to have an exact solution: $u(t)=\frac{1}{2}+\frac{1}{2}tanh\left(\frac{5}{8}t-\frac{x}{4}\right)$
We report in Figure 19 the results obtained. It is important to note that the red dots on the ”bottom” of the domain are points in which the initial condition is enforced, while on the two sides the boundary conditions are enforced as in the previous examples.
The nonlinearity has been solved by Gauss-Newton iterations, with a control on the norm of the update. This method converges in few iterations and kepps the accuracy that was obtained in the linear cases.
Figure 20 shows the convergence of the error on test points as the number of training points and neurons in the hidden layer increase. Notably, the convergence rate for this nonlinear example is similar to other linear problems with similar regularity in the solution.
Figure 19: Example for the over-parametrized network with $N=1000$. On the left the distribution of training points, on the right the computed error on logarithmic scale
Figure 20: Convergence plot for the square (same number of neurons as training points) and over-parametrized cases.
5.7 2D Nonlinear Diffusion
In the last example, we consider a nonlinear diffusion: $\frac{\partial u}{\partial t}=u\Delta u$ The spatial domain is $\Omega=[0,1]^{2}$ and the equation is solved for values of $t$ in the interval $[0,0.25]$. The Dirichlet boundary conditions are fixed to have an exact solution: $u=\frac{1}{1-t}\left(\frac{x^{2}+y^{2}+1}{4}+x\,y+x+y\right)$
We report in Figure 21 the evolution in time of the spatial error obtained. A little increment as time goes on is expected since the gradient of the solution also increases
Figure 22 shows the convergence of the error on test points as the number of training points and neurons in the hidden layer increase. The method is able to achieve good accuracy with a similar amount of training points, despite the fact that we are working on a 3D domain. From the figure, it can be noticed that the algorithm starts to converge when a sufficient number of points is taken and reaches a good accuracy rapidly, while it cannot gain higher accuracy when increasing the number of degrees of freedom.
Figure 21: The error is computed on the spatial domain at each instant, then the values are plotted as a function of time
Figure 22: Convergence plot for the square and over-parametrized cases.
@@ 6 Conclusions
In the present work, we have presented a new method for the numerical resolution of linear and nonlinear differential problems using single hidden-layer neural networks. This strategy is able to consider boundary conditions in an efficient way and mitigates some negative aspects of the previous strategies. The proposed LSE-ELM method is a global (space-time) collocation strategy, and this gives many pros and some cons. The pros include the possibility to consider points outside the domain in the resolution, so as to move the computation in the direction of physically based extrapolation. In fact, in the resolution of PDEs from the point of view of training, one can consider that the usual boundary conditions are good ”supervised data”, while the satisfaction of the differential problem can be a ”physical constraint”. Then, the resolution both inside the domain and outside can be an important feature in applications where extrapolation is required. The cons include the resolution of a full matrix and the loss of accuracy when the problem is less regular. The latter is an issue that can be fixed only with local strategies or the use of ad hoc activation functions. For the first contra, we have noticed with some computations that the spectra are better than the direct competitors, and we believe that some adaptation can further improve such a task. The constructions in the present paper are done in the most general case, and no adaptivity has been included, as an example with the use of Null Rules [2, 3] one could adapt the shape of the functions. However, the strategy has been used in such a setting in order to emphasize the ability of the method to deal with randomly chosen points and activation functions.